OpenStudy (anonymous):

Find the leg of each isosceles right triangle when the hypotenuse is of the given measure. Given = 8

OpenStudy (anonymous):

assume each leg = x then using pythagoras theorem, x^2 + x^2 = (hypotenuse)^2 2x^2 = 8^2 2x^2 = 64 x^2 = 32 x = sqrt(32) x = 4* sqrt(2)

OpenStudy (anonymous):

thats what i got but it says that its wrong

OpenStudy (anonymous):

how can that be? have they given u the answer?